# Generic motives and motivic cohomology of fields

### v1

@@ 2.1. Notation from motivic homotopy
The construction of the category $\operatorname{DM}(k)$ via sheaves with transfers has been extended to arbitrary noetherian finite dimensional bases in [CD19]. One should be careful that over a positive dimensional base, the $\mathbb{A}^{1}$-local complexes (which form the effective category, as above) do not admit a simple description as above. They are given by complexes whose (hyper)cohomology is $\mathbb{A}^{1}$-invariant. One can define a stable category, using essentially the same procedure as above. It admits all of the six functors but, up to now, the localization property for closed immersion has only been proved in some restricted cases (see [CD19, 11.4] for details).
It was shown in [CD15] that, up to replacing the Nisnevich topology with the $\mathrm{cdh}$-topology, working over $k$-schemes $S$ for a fixed base field $k$, and inverting the characteristic exponent of $k$, then localization holds and therefore, all of the six functors formalism (including duality) is available for these “$\mathrm{cdh}$-motives”. Besides, they coincide with “modules over motivic cohomology”, that is more precisely, motives over the motivic Eilenberg-MacLane spectrum as defined by Voevodsky.
The strategy of defining motives as modules over an appropriate spectrum was then developed by Markus Spitzweck in [Spi18], where he defines an apropriate Eilenberg-MacLane motivic spectrum over a Dedekind scheme (say $\mathbb{Z}$) using Bloch’s higher Chow groups. Taking pullbacks in the stable motivic homotopy category then provides the correct $E_{\infty}$-ring spectrum. Finally, modules over this $E_{\infty}$-spectrum inherits all of the properties of the 6 functors formalism from motivic stable homotopy — this last step was already highlighted in the abstract case and in the rational coefficients case in [CD19, §7 and Th. 14.2.9] so that we know that Spitzweck’s construction and the one highlighted above for the Nisnevich topology agrees with rational coefficients over geometrically unibranch schemes. For arbitrary singular schemes and rational coefficients, this also holds provided one uses the $\mathrm{cdh}$-topology.
2.2. Motivic cohomology
@@
It follows from the so-called continuity property (see [CD19, 11.1.24, 11.1.25]) that these cohomology groups commute with projective limits of schemes whenever the transition morphisms of an essentially affine pro-scheme $(X_{\lambda})_{\lambda\in\Lambda}$ are dominant (see loc. cit. for recall). So when $X=\lim_{\lambda\in\Lambda}X_{\lambda}$ we get under the preceding assumptions:
@@
These motivic cohomology groups can be computed with the Nisnevich cohomology of the $i$-th motivic complex $\mathbbm{1}(i)=\mathbb{Z}(i)=\operatorname{C_{*}^{sus}}(\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i}))$: [^12] Here $\mathbb{G}_{m}^{\wedge,i}$ means the smash product of the pointed scheme $(\mathbb{G}_{m},1)$, and $\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i})$ denotes the associated sheaf with transfers. Equivalently, $\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i})$ is the cokernel of the map: $\oplus_{i=1}^{n}\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{i-1})\rightarrow\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{n})$ obtained by considering all the possible inclusions.
@@
This isomorphism follows from Voevodsky’s theory (recalled above) when $k$ is perfect. In the general case, $k$ is essentially smooth over its prime field $F$, so that $X$ is also essentially smooth over $F$, and one reduces to the case of $F$ using the continuity property. Along with the (quasi-)isomorphisms of motivic complexes
@@
this implies the following computations of these motivic cohomology groups:
@@
with respectively is the classical Chow groups of codimension $n$ algebraic cycles of $X$ modulo rational equivalence, and the $n$-th Milnor K-theory of any function field $E/k$.
Recall finally that it is known due to a theorem of Voevodsky that motivic cohomology coincides with Bloch’s higher Chow groups, [^13] Bloch’s definition, motivated by earlier results of Kratzer and Soulé, and by the ongoing excitement that emerged from the picture laid down by Lichtenbaum and Beilinson, appeared around 10 years before Voevodsky’s definition. For the record, it was the will of repairing a mistake in the proof of the localization property for higher Chow groups that motivated Suslin to introduce Suslin homology in a talk in the CIRM at Luminy in 1987. This seminal definition was a major contribution that shaped Voevodsky’s theory of motivic complexes a few years later. Meanwhile, the mistake was fixed by Bloch [Blo94] and Levine [Lev94]. according to the following formula:
@@
Rational coefficients. Recall that the rational part of the above motivic cohomology groups can be compared with Quillen higher algebraic K-theory according to the following fundamental formula:
@@
where we have denoted by $K_{2i-n}^{(i)}(X)$ the $i$-th graded parts for the $\gamma$-filtration on rational K-theory $K_{2i-n}(X)_{\mathbb{Q}}$. In [Sou85], the right hand-side was actually taken as a definition of motivic cohomology, following the lead of Beilinson. For this reason, we also use the notation (due to Riou):
@@
In the case of a smooth $k$-scheme, the above formula is a consequence Voevodsky’s isomorphism (2.2.1.e), and the corresponding isomorphism for rational higher Chow groups due to Bloch ([Blo86]). A more general proof, valid for an arbitrary regular scheme $X$ (without a base field), is given in [CD19, 14.2.14, 16.1.4]. [^14] For singular schemes $X$, several works have attempted to propose a suitable definition for an integral version of the left hand-side of (2.2.2.a), whose rational part coincides with Soulé’s definition. This is still an active area of research.
2.3. Pro-object
Our references for pro-objects is [AM69, Appendix] for $1$-categories, and [Lur09, §5.3], [BHH17] for $\infty$-categories.
Let $\mathcal{C}$ be a presentable $\infty$-category. The $\infty$-category $\operatorname{pro-\mskip-3.0mu}\mathcal{C}$ admits cofiltered (homotopy) projective limits. In particular, given an $\infty$-functor $X:\mathscr{I}\rightarrow\mathcal{C},i\mapsto X_{i}$, one can consider the associated projective limit in $\operatorname{pro-\mskip-3.0mu}\mathcal{C}$ that we will denote by “$\underset{i\in\mathscr{I}}{\text{\rm``lim''}}X_{i}$ following a convention of Deligne.
According to [BHH17, Prop. 5.1.1, 5.1.2], given two $\infty$-functors $X:\mathscr{I}\rightarrow\mathcal{C}$ and $Y:\mathscr{J}\rightarrow\mathcal{C}$, there exists a canonical equivalence:
$\operatorname{Map}(\underset{i\in\mathscr{I}}{\text{\rm``lim''}}X_{i},\underset{j\in\mathscr{J}}{\text{\rm``lim''}}Y_{j})\simeq\lim_{j\in\mathscr{J}}\underset{i\in\mathscr{I}}{\operatorname{colim}}\operatorname{Map}(X_{i},Y_{j}).$
One should be careful that in general: $\operatorname{Ho}(\operatorname{pro-\mskip-3.0mu}\mathcal{C})\neq\operatorname{pro-\mskip-3.0mu}\operatorname{Ho}\mathcal{C}$ (but see 3.1.11).
3. The category of Generic motives
@@ 3.1. Generic motives and points for the homotopy $t$-structures
We will use the following handy definition.
@@
A $k$-scheme $\mathcal{X}$ is pro-smooth (resp. essentially smooth) if it is the projective limit of a pro-scheme $(X_{i})_{i\in I}$ such that $X_{i}$ is a $k$-scheme of finite type, and the transition maps are étale (resp. open immersions) and affine.
According to [EGA4, 8.13.2], the category of pro-smooth (resp. essentially smooth) $k$-schemes can be viewed as a full sub-category of $\operatorname{pro-\mskip-3.0mu}\mathrm{Sm}$. In fact, the functor which to a pro-scheme $(X_{i})_{i\in I}$ as above associate the projective limit $\mathcal{X}=\lim_{i\in I}X_{i}$ seen as an object of $\overline{\mathrm{Sm}}$ is an equivalence of categories. We will say that $(X_{i})_{i\in I}$ is a presentation of $\mathcal{X}$, keeping in mind the choice of a particular presentation is irrelevant. In particular, we will consider the embedding [^15] Of course, it formally corresponds to the choice of a reciprocal equivalence to the projective limit functor:
(3.1.1.a) $\nu:\overline{\mathrm{Sm}}\rightarrow\operatorname{pro-\mskip-3.0mu}\mathrm{Sm}_{k}.$
@@
    As in [Dég08b], one can use a canonical presentation of $\operatorname{Spec}(E)$, by considering the set $\mathcal{M}(E)$ of smooth sub-$k$-algebras $A\subset E$ with fraction field equal to $E$. Ordered by inclusion, this set is filtered and one has:
    $\operatorname{Spec}(E)=\lim_{A\in\mathcal{M}(E)^{op}}\operatorname{Spec}(A).$
    One gets another presentation by considering a smooth model [^16] Resolution of singularities leads to the existence of a smooth projective model. Therefore, such a model exists in characteristic 00 or when the transcendence degree of $E/k$ is less than $3$. of $E$: that is a connected smooth $k$-scheme $X$ with a chosen $k$-isomorphism $\phi:E\simeq\kappa(X)$ Then one gets a canonical isomorphism:
    $\operatorname{Spec}(E)=\lim_{U\subset X}U$
    where $U$ runs over the non-empty open subset of $X$.
@@
The composite functor:
$\mathrm{Sm}_{k}\xrightarrow{\operatorname{M}}\operatorname{DM_{gm}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k),X\mapsto\operatorname{M}(X)$
admits a right Kan extension $\operatorname{\hat{M}}$ along the inclusion $\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$, in the $\infty$-categorical sense ([Cis19, §6.4]). This formally follows from the fact $\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)$ is complete171717i.e. admits all small limits; which follows by construction as an $\infty$-category, but one gets a direct construction by using the following commutative diagram:
@@
In other words,
@@
for any presentation $(X_{i})_{i\in I}$ of the essentially smooth $k$-scheme $\mathcal{X}$. As in [Dég08b, Def. 3.3.1], one highlights the following pro-motives:
@@
The $\infty$-category $\operatorname{DM^{(0)}}(k)$ of generic motives (over $k$) is the full sub-$\infty$-category of $\operatorname{pro-\mskip-3.0mu}\operatorname{DM_{gm}}(k)$ whose objects are of the form $\operatorname{\hat{M}}(E)\{n\}$ for any function field $E/k$ and any integer $n\in\mathbb{Z}$.
@@
One should be careful that the above definition of generic motives is sensibly different to that of loc. cit. as it includes the higher homotopy information, contained in the mapping spaces of the $\infty$-category $\operatorname{DM^{(0)}}(k)$. However, we will see in 3.1.12 that the homotopy category of the later coincides with the category defined in loc. cit.
@@
The reason to focus on these particular pro-objects is that they form a conservative family of points for the $t$-category $\operatorname{DM}(k)$ equipped with its homotopy $t$-structure. Before giving a precise meaning to this sentence in the following corollary, we introduce another notation. Given an $R$-linear presheaf $F$ on $\mathrm{Sm}_{k}$, we denote by $\hat{F}$ its left Kan extension along the inclusion $\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$. Concretely, for any essentially smooth $k$-scheme $\mathcal{X}$, we have:
$\hat{F}(\mathcal{X})=\operatorname{colim}_{i\in I}F(X_{i})$
for any presentation $(X_{i})_{i\in I}$ of $\mathcal{X}$.
@@
Let $\mathcal{O}$ be an essentially smooth henselian semi-local $k$-algebra, $(n,i)\in\mathbb{Z}^{2}$ a pair of integers, and $\mathbf{E}$ be a stable motivic complex over $k$. Then there exists a canonical isomorphism:
$\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(\operatorname{\hat{M}}(\mathcal{O})\{i\}[n],\mathbf{E})=\hat{\operatorname{\underline{H}}}_{n,-i}(\mathbf{E})(\mathcal{O})$
where $\operatorname{\underline{H}}_{n,*}(\mathbf{E})$ is the $n$-th homology object associated with $\mathbf{E}$, seen as a homotopy module. [^18] So the symbol * refers to the natural graduation of homotopy modules, which actually coincides with the $\mathbb{G}_{m}$-graduation on $\mathbf{E}$.
@@
First recall that $\operatorname{\underline{H}}_{n,*}(\mathbf{E}\{-i\})=\operatorname{\underline{H}}_{n,*-i}(\mathbf{E})$. This reduces to the case $i=0$. Then, using the cancellation theorem (recalled in 2.1.2), one can compute the relevant group morphism in $\operatorname{DM^{eff}}(k)$. Then one uses the hypercohomology spectral sequence for the homotopy $t$-structure and one concludes as $\mathcal{O}$ is acyclic for Nisnevich cohomology with coefficients in an object of $\operatorname{HI}(k)$ (see e.g. [Ros96, Th. 6.1] taking into account [Dég11, 3.10, 3.11]). ∎
@@
Let $X$ be a smooth projective $k$-scheme of dimension $d$. Using duality (see 2.1.3) and the continuity property (2.2.1.a), one immediately deduces the following computation:
$\hat{\operatorname{\underline{H}}}_{n,-i}(\operatorname{M}(X))(E)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{DM}(k)}(\operatorname{\hat{M}}(E)\{i\}[n],\operatorname{M}(X))=\operatorname{H}_{\mathrm{M}}^{2d-n-i,d-i}(X_{E})$
@@
One deduces the announced interpretation of generic motives as a conservative family of points with respect to the homotopy $t$-structure. [^19] In fact, a much stronger property holds: see 3.3.7 and 3.3.8.

### v3

@@ 2.1. The $\infty$-category of motivic complexes and its stabilization
The construction of motivic complexes, effective and stable, has been extended to arbitrary noetherian finite dimensional bases $S$. This is based on a strategy started by Voevodsky in [Voe10b], using the extension of finite correspondences due to Suslin and Voevodsky. The explicit construction has been written down in [CD19, Part III], which gives appropriate monoidal model categories, and therefore presentable stable monoidal $\infty$-categories $\operatorname{\mathscr{D}M^{eff}}(S)$ and $\operatorname{\mathscr{D}M}(S)$. It follows the same pattern as above, starting from the derived $\infty$-category $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(S))$ of sheaves with transfers over $S$. As an example, the objects of $\operatorname{DM^{eff}}(S)$ can similarly be described as $\mathbb{A}^{1}$-local complexes in $\operatorname{\mathscr{D}}(\operatorname{Sh^{tr}}(S))$. [^18] One should be careful that the simpler description of $\mathbb{A}^{1}$-local objects obtained over a perfect field using 2.1.2 fails over positive dimensional bases, as shown by Joseph Ayoub in [Ayo06]. At this time, it is known that $\operatorname{\mathscr{D}M}(S)$ admits all of the six functors, but the localization property for closed immersions has only been established in certain restricted cases (see [CD19, 11.4] for details).
On the other hand, an alternative strategy to define an appropriate category of stable motivic complexes was outlined by Oliver Röndigs and Paul Arne Østvær in their work [RØ08]. Indeed, in loc. cit. they proved that for a field $k$ of characteristic 00, the triangulated category $\operatorname{DM}(k)$ is equivalent to the category of modules over the Eilenberg-MacLane motivic spectrum defined by Voevodsky ([Voe02b]). Since the latter object was also defined over an arbitrary base, one could try to look at modules over some arbitrary base $S$ to define motivic complexes over $S$. This strategy can be made effective, but one still runs into a major problem in order to get the six functors formalism: one has to prove that the motivic Eilenberg MacLane is stable under pullbacks, which is the last conjecture, conjecture 17, of the paper [Voe02b].
This conjecture was solved in [CD19, 16.1.7] for rational coefficients and a geometrically unibranch base scheme $S$. More generally, one can show that the rationalization of $\operatorname{\mathscr{D}M}(S)$ is equivalent to the category of modules over an appropriately defined rational motivic Eilenberg-MacLane spectrum ([CD19, 14.2.9]). Integrally, Voevodsky’s conjecture is known for pullbacks of regular $k$-schemes where $k$ is a fixed base field; see [CD15, Cor. 3.6]. This result enabled the authors to realize the program of Röndigs and Østvær in their work. Furthermore, by restricting attention to $k$-schemes, inverting the characteristic exponent of $k$, and replacing the Nisnevich topology with the $\mathrm{cdh}$-topology, it is proved in loc. cit. that $\mathrm{cdh}$-local motivic complexes do satisfy the localization property, are equipped with the six functors formalism, and are equivalent to the category of modules over the appropriately defined Eilenberg-MacLane motivic ring spectrum.
Finally, the most successful strategy to get a working theory of stable motivic complexes over arbitrary bases has been developed by Markus Spitzweck in [Spi18], still adopting the general strategy of Röndigs and Østvær. Drawing on Bloch’s higher Chow groups and their extension over Dedekind domains by Marc Levine, Spitzweck constructs a suitable Eilenberg-MacLane motivic spectrum over a Dedekind scheme. By taking pullbacks in the stable motivic homotopy category, one then obtains the desired $E_{\infty}$-ring spectrum.
Modules over this $E_{\infty}$-ring spectrum inherit all the expected properties of the six functors formalism from the motivic stable homotopy category as established by Ayoub in [Ayo07b]. With rational coefficients, Spitzweck’s theory coincides with that of [CD19] and that of [Ayo14].
2.2. A quick review of motivic cohomology
@@
It follows from the so-called continuity property (see [CD19, 11.1.24, 11.1.25]) that these cohomology groups commute with projective limits of schemes whenever the transition morphisms of an essentially affine pro-scheme $(X_{\lambda})_{\lambda\in\Lambda}$ are dominant (see loc. cit. for details). So when $X=\operatorname{lim}_{\lambda\in\Lambda}X_{\lambda}$ we get under the preceding assumptions:
@@
These motivic cohomology groups can be computed with the Nisnevich cohomology of the $i$-th motivic complex $\mathbbm{1}(i)=\mathbb{Z}(i)=\operatorname{C_{*}^{sus}}(\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i}))$: [^19] Here, $\mathbb{G}_{m}^{\wedge,i}$ denotes the smash product of the pointed scheme $(\mathbb{G}_{m},1)$, and $\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i})$ denotes the associated sheaf with transfers. Equivalently, $\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{\wedge,i})$ is the cokernel of the map: $\oplus_{i=1}^{n}\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{i-1})\rightarrow\operatorname{\mathbb{Z}^{tr}}(\mathbb{G}_{m}^{n})$ obtained by considering all the possible inclusions.
@@
This isomorphism follows from Voevodsky’s theory (recalled above) when $k$ is perfect. In the general case, $k$ is essentially smooth over its prime field $F$, so $X$ is also essentially smooth over $F$, and one to reduce to the case of $F$ using the continuity property. Along with the (quasi-)isomorphisms of motivic complexes
@@
this leads to the following computations of the motivic cohomology groups:
@@
with respectively the classical Chow groups of codimension $n$ algebraic cycles of $X$ modulo rational equivalence, and the $n$-th Milnor K-theory of any function field $E/k$.
Recall finally that it is known due to a theorem of Voevodsky ([Voe02a]) that motivic cohomology coincides with Bloch’s higher Chow groups, [^20] Bloch’s definition, motivated by earlier results of Kratzer and Soulé, and by the ongoing excitement that emerged from the picture laid down by Lichtenbaum and Beilinson, appeared around 10 years before Voevodsky’s definition. For the record, it was the desire to correct a mistake in the proof of the localization property for higher Chow groups that motivated Suslin to introduce Suslin homology in a talk in the CIRM at Luminy in 1987. This seminal definition was a major contribution that shaped Voevodsky’s theory of motivic complexes a few years later. The error was subsequently fixed by Bloch [Blo94] and Levine [Lev94]. according to the following formula:
@@
Rational coefficients. Recall that the rational part of the above motivic cohomology groups can be compared with Quillen’s higher algebraic K-theory according to the following fundamental formula:
@@
where we have denoted by $K_{2i-n}^{(i)}(X)$ the $i$-th graded parts for the $\gamma$-filtration on rational K-theory $K_{2i-n}(X)_{\mathbb{Q}}$. As recalled in the introduction, the right-hand side was actually taken as a definition of rational motivic cohomology in [Sou85] and [Bei84]. For this reason, we also use the notation due to Riou:
@@
In the case of a smooth $k$-scheme, the above formula is a consequence of Voevodsky’s isomorphism (2.2.1.e), and the corresponding isomorphism for rational higher Chow groups due to Bloch ([Blo86]). A more general proof, valid for an arbitrary regular scheme $X$ (without a base field), is given in [CD19, 14.2.14, 16.1.4]. [^21] For singular schemes $X$, several authors have proposed suitable definitions for an integral version of the left-hand side of (2.2.2.a), whose rational part coincides with Soulé’s definition. This is still an active area of research.
3. Review and complement on generic motives
@@ 3.1. Generic motives and points for the homotopy $t$-structures
We will use the following convenient definition.
@@
A $k$-scheme $\mathcal{X}$ is said to be pro-smooth (resp. essentially smooth) if it is the projective limit of a pro-scheme $(X_{i})_{i\in I}$ such that $X_{i}$ is a $k$-scheme of finite type, and the transition maps are étale (resp. open immersions) and affine.
We let $\overline{\mathrm{Sm}}_{k}$ be the category of pro-smooth $k$-schemes.
According to [GD67, 8.13.2], the category of pro-smooth (resp. essentially smooth) $k$-schemes can be viewed as a full subcategory of $\operatorname{pro-\mskip-3.0mu}\mathrm{Sm}$. In fact, the functor $P$ that assigns to a pro-scheme $(X_{i})_{i\in I}$ as above the projective limit $\mathcal{X}=\operatorname{lim}_{i\in I}X_{i}$ seen as an object of $\overline{\mathrm{Sm}}_{k}$ is an equivalence of categories. We will say that $(X_{i})_{i\in I}$ is a presentation of $\mathcal{X}$, keeping in mind that the choice of a specific presentation is irrelevant. To formulate the next results, we consider an arbitrary quasi-inverse to the functor $P$:
(3.1.1.a) $\nu:\overline{\mathrm{Sm}}_{k}\rightarrow\operatorname{pro-\mskip-3.0mu}\mathrm{Sm}_{k}.$
@@
    As in [Dég08b], one can use a canonical presentation of $\operatorname{Spec}(E)$ by considering the set $\mathcal{M}(E)$ of smooth sub-$k$-algebras $A\subset E$ whose fraction field is $E$. Ordered by inclusion, this set is filtered and one has:
    $\operatorname{Spec}(E)=\operatorname{lim}_{A\in\mathcal{M}(E)^{op}}\operatorname{Spec}(A).$
    One gets another presentation by considering a smooth model of $E$: that is a connected smooth $k$-scheme $X$ with a chosen $k$-isomorphism $\phi:E\simeq\kappa(X)$ This leads to a canonical isomorphism:
    $\operatorname{Spec}(E)=\operatorname{lim}_{U\subset X}U$
    where $U$ ranges over the non-empty open subsets of $X$.
@@
The composite $\infty$-functor:
$\mathrm{Sm}_{k}\xrightarrow{\operatorname{M}}\operatorname{\mathscr{D}M_{gm}}(k)\rightarrow\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k),\quad X\mapsto\operatorname{M}(X)$
admits a right Kan extension $\operatorname{\hat{M}}$ along the inclusion $\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$, in the $\infty$-categorical sense (see [Cis19, §6.4]). This follows formally from the fact $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)$ is complete222222i.e. admits all small limits, which follows by construction as an $\infty$-category, but one gets a direct construction by using the following commutative diagram:
@@
Indeed, following our convention for pro-objects (see 6.6.4), we have the following explicit formula:
@@
for any presentation $(X_{i})_{i\in I}$ of the essentially smooth $k$-scheme $\mathcal{X}$. Following [Dég08b, Def. 3.3.1], we highlight the following pro-motives:
@@
The $\infty$-category $\operatorname{\mathscr{D}M^{(0)}}(k)$ of generic motives (over $k$) is defined as the full sub-$\infty$-category of $\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M_{gm}}(k)$ whose objects are of the form $\operatorname{\hat{M}}(E)\{n\}$ for any function field $E/k$ and any integer $n\in\mathbb{Z}$.
@@
One should be mindful that the above definition of generic motives is substantially different from that in loc. cit. as it retains the higher homotopy information, contained in the mapping spaces of the $\infty$-category $\operatorname{\mathscr{D}M^{(0)}}(k)$. However, we will see in 3.2.3 that the homotopy category of the latter coincides with the category introduced in loc. cit.
@@
The reason for focusing on these particular pro-objects is that they form a conservative family of points for the $t$$\infty$-category $\operatorname{\mathscr{D}M}(k)$, equipped with its homotopy $t$-structure. Before making this assertion precise in the following corollary, we introduce a new notation: given an $R$-linear presheaf $F$ on $\mathrm{Sm}_{k}$, we denote by $\hat{F}$ its left Kan extension along the inclusion $\mathrm{Sm}_{k}\rightarrow\overline{\mathrm{Sm}}_{k}$. Concretely, for any essentially smooth $k$-scheme $\mathcal{X}$, we have:
$\hat{F}(\mathcal{X})\simeq\operatorname{colim}_{i\in I}F(X_{i})$
for any presentation $(X_{i})_{i\in I}$ of $\mathcal{X}$. When $F_{*}$ is a homotopy module (see 2.1.4), we simply denote by $\hat{F}_{*}$ the extension to $\overline{\mathrm{Sm}}$ just defined. In particular, when $\mathbf{E}$ is a motivic spectrum, with associated $n$-th homology object $\operatorname{\underline{H}}_{n,*}(\mathbf{E})$ seen as a homotopy module (see again 2.1.4), we will denote by $\hat{\operatorname{\underline{H}}}_{n,*}(\mathbf{E})$ the extension to $\overline{\mathrm{Sm}}$ defined above.
@@
Let $\mathcal{O}$ be an essentially smooth semi-local $k$-algebra, $(n,i)\in\mathbb{Z}^{2}$ a pair of integers, and $\mathbf{E}$ a stable motivic complex over $k$. Then there exists a canonical isomorphism:
$\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(\mathcal{O})\{i\}[n],\mathbf{E})\simeq\hat{\operatorname{\underline{H}}}_{n,-i}(\mathbf{E})(\mathcal{O})$
where on the left-hand side, $\operatorname{\hat{M}}(\mathcal{O})=\operatorname{\hat{M}}(\operatorname{Spec}(\mathcal{O}))$ and $\mathbf{E}$ is seen as a constant pro-object.
@@
First, recall that $\operatorname{\underline{H}}_{n,*}(\mathbf{E}\{-i\})=\operatorname{\underline{H}}_{n,*-i}(\mathbf{E})$. This reduces to the case $i=0$. Next, using the cancellation theorem (recalled in 2.1.4), one can compute the relevant group morphism in $\operatorname{\mathscr{D}M^{eff}}(k)$. We then apply the hypercohomology spectral sequence for the homotopy $t$-structure, extended to the essentially smooth $k$-scheme $\mathcal{X}=\operatorname{Spec}(\mathcal{O})$, which has the following form:
$E_{2}^{p,q}=H^{p}(\mathcal{X},\operatorname{\underline{H}}_{-q,0}(\mathbf{E}))\Rightarrow\mathbf{E}^{p+q}(\mathcal{X}).$
The $E_{2}$-term is the cohomology with coefficients in a homotopy module, and this coincides with the Chow groups with coefficients in the associated cycle module, according to[Dég11, 3.10, 3.11]. One deduces from the Gersten conjecture for Chow groups with coefficients, as proved in [Ros96, Th. 6.1], that the $E_{2}$-term of the above spectral sequence is concentrated on the line $p=0$. This concludes the proof. ∎
@@
Let $X$ be a smooth projective $k$-scheme of dimension $d$. Using duality (see 2.1.5) and the continuity property (2.2.1.a), one immediately deduces the following computation:
$\hat{\operatorname{\underline{H}}}_{n,-i}(\operatorname{M}(X))(E)=\operatorname{Hom}_{\operatorname{pro-\mskip-3.0mu}\operatorname{\mathscr{D}M}(k)}(\operatorname{\hat{M}}(E)\{i\}[n],\operatorname{M}(X))=\operatorname{H}_{\mathrm{M}}^{2d-n-i,d-i}(X_{E})$
@@
3.1.9.
We now turn to the promised interpretation of generic motives as “points” for the homotopy $t$-structure on $\operatorname{\mathscr{D}M}(k)$. The term points is used here by analogy with the more classical notion of points in topos theory. Indeed, a point of a topos can be seen as a fiber functor, an exact functor to the punctual topos.
In our setting, we will leverage the fact (see 6.4.5) that any stable $\infty$-category admits mapping spaces with values in the $\infty$-category of spectra $\mathscr{S}p$, which we see as the analog of the punctual topos. [^23] In fact, the $\infty$-category $\mathscr{S}p$ is the initial presentable stable $\infty$-category. Then the exactness property of fiber functors is replaced by the $t$-exactness property, with respect to the standard (Postnikov) $t$-structure on spectra.